OpenStudy (anonymous):

How do i factor (a+3)^2

geerky42 (geerky42):

It is already factored?

OpenStudy (anonymous):

it is?

OpenStudy (imtiaz7):

likesquare formula: (a+b)^2= a^2+2ab+b^2 and here a=a and b=3

OpenStudy (anonymous):

huh

geerky42 (geerky42):

yeah it is already factored, or maybe you just square it and have \((a+3)(a+3)\). I think you mean " expand \((a+3)^2\) "

OpenStudy (anonymous):

it doesnt make sense to expand it.

OpenStudy (anonymous):

actually youre right. Thanks